OpenStudy (anonymous):

what is the value of z so that -9 and 9 are both solutions of x^2 + z = 103? A)-22 B) 3 C) 22 D) 184

OpenStudy (anonymous):

can you help me @rose21

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

@kewlgeek555

OpenStudy (rose21):

maybe @mathmale or @kewlgeek555 might be able to help

OpenStudy (jack1):

x^2 + z = 103 9^2 + z = 103 81 + z = 103 z = 103 - 81 = ...?

OpenStudy (jack1):

same answer for x = -9 as -9 * -9 also = 81

OpenStudy (anonymous):

so it is 22?

OpenStudy (kewlgeek555):

I think Mr. Jack the First provides an excellent explanation. :3 Thanks for calling me though. ;D

OpenStudy (anonymous):

Thank you @kewlgeek555

OpenStudy (anonymous):

and thank you for the help @Jack1

OpenStudy (jack1):

hall good dude, yw hey